1/xy2+y4 is an integral factor of the differential equation(x2y+y2)dx+(y3-z3)dy=0 true or false


1
Expert's answer
2021-04-20T15:49:00-0400

Given equation,

"(x^2y+y^2)dx+(y^3-x^3)dy=0"


Here, "M=x^2y+y^2, N=y^3-x^3"


Hence Integrating factor "I.F.=\dfrac{1}{Mx+Ny}=\dfrac{1}{x^3y+xy^2+y^4-x^3y}"


"=\dfrac{1}{xy^2+y^4}"


Hence "\dfrac{1}{xy^2+y^4}" is an integrating factor of given equation.